# Beyond Formal Semantics for Capabilities and Skills: Model Context Protocol in Manufacturing

### v1

@@ Abstract
Index Terms:
Model Context Protocol, MCP, Capabilities, Skills, Large Language Models, LLMs, Ontologies, Semantic Web
@@ I Introduction
Modern industrial automation systems are increasingly expected to operate in dynamic environments, where production requirements, tasks, and system configurations change frequently. To meet these demands, such systems must be both modular and reconfigurable [1].
A common approach to describe and manage the functional potential of these systems is the use of Capabilities, Skills and Services (CSS). Capabilities describe abstract, implementation-independent functions, whereas skills represent their concrete implementations and are exposed via a skill interface [2]. This separation supports abstraction, reuse, and planning by allowing reasoning on the capability level while delegating execution to the skill level. However, the practical use of capabilities and skills in flexible automation faces a persistent challenge in the formal modeling of capabilities and skills. This model creation is labor-intensive, error-prone, and demands deep semantic modeling expertise [3].
At the same time, Large Language Models (LLMs) have emerged as powerful reasoning tools capable of analyzing complex problems, generating plans, and supporting decision-making processes. However, LLMs do not natively have access to functionalities like machine interfaces, control systems, or sensor values. To integrate external functionalities with LLMs, Model Context Protocol (MCP) was recently introduced [4]. LLMs can interpret task descriptions, reason about available functionalities, select and invoke suitable functions via MCP, and thus contribute to flexible planning and control.
@@ II Model Context Protocol
Model Context Protocol (MCP), introduced in [4], is an open standard for connecting LLMs to external resources, prompt templates and tools, i.e., executable functions, in a structured and unified way. While an LLM alone may conceptually plan a task for a robotic cell, it cannot efficiently execute the corresponding low-level operations without a standardized mechanism to invoke such external functions. With MCP, a protocol layer is provided that allows external functions to be described, discovered, and invoked at runtime – without requiring prior hard-coded integration into the LLM.
MCP follows a client-server architecture: MCP servers expose functional interfaces, metadata, and invocation mechanisms; MCP clients, e.g., LLM-based agents, can dynamically query these servers to discover available operations and execute them by issuing well-defined requests. The communication between MCP clients and servers is based on JSON-RPC 2.0, allowing functions to be called with structured parameters and standardized response handling. [4]
@@ III Related Work
A recent survey in [5] outlines MCP ’s core concepts and its potential to expose external tools to LLMs in a standardized way. The authors discuss possible applications of MCP in domains such as manufacturing, where LLM agents might retrieve sensor data, interact with MES platforms, or initiate control actions in response to anomalies. Ref. [5] highlights opportunities for using MCP in planning and control within smart factories, but it does not present concrete implementations or evaluate the feasibility of using LLMs for tool selection or task execution in physical automation systems.
@@
While the authors of [5] outline the conceptual potential of the new MCP for exposing tools to LLMs, other recent studies focus more directly on how LLMs can be applied to planning and control tasks in industrial automation. Xia et al. present in [6] an agent-based framework for LLM-controlled industrial automation. The presented framework relies on a structured prompting scheme, event-driven semantics, and a hierarchy of manager and operator agents that interact with automation modules via digital twins. The control is carried out by generated function calls based on current event logs, with standard operation procedures serving as a behavior-guiding knowledge base. However, this approach requires fine-tuning, curated datasets, and tightly engineered prompt structures – limiting flexibility and generalizability.
@@
A similar direction is pursued in the LLMAPM framework introduced in [7], which targets manufacturing process planning in Industry 5.0. Natural language task descriptions are decomposed into workflows, verified through state machines, and executed on heterogeneous devices via a low-code platform. While supporting human-in-the-loop corrections, the system still depends on handcrafted prompt templates and lacks a runtime interface for tool abstraction and reuse.
@@
In [8] a multi-agent LLM control framework for robust decision-making in industrial settings is introduced. Their architecture uses validator and reprompter agents with a digital twin to iteratively refine actions. While this enables safe and adaptive control, it is demonstrated on a single closed-loop task and does not generalize to dynamic function composition or modular tool invocation.
@@
The use of LLMs in multi-robot systems is reviewed by Li et al. in [9], covering applications in task allocation and coordination. While relevant for distributed systems, the focus remains on agent negotiation and motion planning rather than exploring how LLMs might interface with standardized protocols like MCP for structured tool invocation.
@@
A broader perspective on LLM-driven agent systems is provided in [10], which categorizes LLM-powered agents into software-based, physical, and adaptive hybrid types. While the survey highlights potential use cases in smart manufacturing, it acknowledges ongoing challenges such as output uncertainty, integration complexity, and lack of standardized tool access in physical environments.
@@ IV-A MCP Architecture
In the proposed architecture (see Figure 1), user interfaces and control systems like MES or SCADA act as MCP clients that are directly connected to an LLM. MCPs servers provide functions, resources, and prompts. According to the definition in [2], a capability is defined as a specification of a function with either physical or virtual effects. This distinction of effects is reflected in the MCP architecture with the two alternatives of direct MCP servers and gateway MCP servers.
@@
User queries or control requests are sent together with all available MCP servers as prompts to the LLM (1a). Depending on the response of the LLM (1b), MCP functionalities are selected and used by the MCP client. In case of direct MCP servers, this functionality is called (2a) and the returned result (2b) is interpreted. In the case of gateway MCP servers, the function call (3a) may lead to an external request (3b) that yields a return (3c), which is processed by the gateway MCP server and then returned to the MCP client (3d). The following subsection demonstrates how elements of the CSS reference model presented in [2] are realized in this MCP architecture.
@@ IV-B Representing CSS Elements in MCP
Each Capability is a function declaration in one of the programming frameworks provided by [4]. Its specification details are documented in natural language using descriptions and code comments. This natural-language description is the counterpart of a capability model and enables LLMs to understand the purpose, constraints, and semantics of a capability.
@@
[…]
A Skill is the implementation of a capability, i.e., the function definition. This is in line with the specification of a skill in [2]. However, MCP provides an alternative interface to expose and describe the skill’s callable entry points.
@@ V Evaluation
    Standard Process Execution: A user specifies a workpiece to be drilled (i.e. defines material and diameter) and transported to an assembly station. The LLM correctly infers that the spindle speed must be calculated first. It uses the calculation skill to retrieve an RPM value, invokes the drilling skill with this RPM value, and dispatches the mobile robot for transport. All skills were executed in correct order and with correct parameters.
@@ VI Conclusion and Future Work
This paper presents a lightweight and LLM-driven alternative to current model-based CSS approaches in manufacturing. By leveraging MCP, we demonstrate how capabilities can be described using comments and skills can be exposed as callable tools. Through a set of scenarios, we show that an LLM paired with MCP can plan and execute multi-step manufacturing processes, discover available capabilities, interpret their purposes, invoke corresponding skills, and handle errors or ambiguities gracefully – without explicit information models, e.g., ontologies or AASs. Our results suggest that explicit CSS models may become less necessary in the future, as LLMs increasingly infer intent from text and function implementations. This would significantly lower the barrier to integration and reconfiguration in flexible manufacturing.
However, our results are only an initial validation of LLM-based orchestration via MCP, and reliability and robustness remain an open challenge with many directions for future work. Instead of a single LLM orchestrator, future approaches should use multiple, dedicated LLM-based agents with defined roles (e.g., planner, verifier, negotiator) for more scalable, modular, and robust architectures. In addition, decentralized negotiation patterns to compare service offers across organizational boundaries may be realized by LLM agents.
Another future research direction is to enable LLMs not only to select and invoke skills, but to also generate new ad-hoc skills on demand. Building on earlier work on skill generation [11], a combined approach could allow LLMs to identify missing functionality, synthesize MCP tool code, register the new tool, and invoke it as part of a task plan. The approach of ”model-free” automation, where planning and execution emerge from linguistic reasoning rather than predefined symbolic models, can also be applied to simplify integration processes in other domains, e.g., modular electrolyzer systems.

### v2

@@ I Introduction
Modern industrial automation systems are increasingly expected to operate in dynamic environments, where production requirements, tasks, and system configurations change frequently. To meet these demands, such systems must be both modular and reconfigurable [ElM_Flexibleandreconfigurablemanufacturing_2005].
A common approach to describe and manage the functional potential of these systems is the use of Capabilities, Skills and Services (CSS). Capabilities describe abstract, implementation-independent functions, whereas skills represent their concrete implementations and are exposed via a skill interface [KBH+_AReferenceModelfor_15.09.2022b]. This separation supports abstraction, reuse, and planning by allowing reasoning on the capability level while delegating execution to the skill level. However, the practical use of capabilities and skills in flexible automation faces a persistent challenge in the formal modeling of capabilities and skills. This model creation is labor-intensive, error-prone, and demands deep semantic modeling expertise [LAF+_TowardaMethodto_2024].
At the same time, Large Language Models (LLMs) have emerged as powerful reasoning tools capable of analyzing complex problems, generating plans, and supporting decision-making processes. However, LLMs do not natively have access to functionalities like machine interfaces, control systems, or sensor values. To integrate external functionalities with LLMs, Model Context Protocol (MCP) was recently introduced [Ant_ModelContextProtocolSpecification_2025]. LLMs can interpret task descriptions, reason about available functionalities, select and invoke suitable functions via MCP, and thus contribute to flexible planning and control.
@@ II Model Context Protocol
Model Context Protocol (MCP), introduced in [Ant_ModelContextProtocolSpecification_2025], is an open standard for connecting LLMs to external resources, prompt templates and tools, i.e., executable functions, in a structured and unified way. While an LLM alone may conceptually plan a task for a robotic cell, it cannot efficiently execute the corresponding low-level operations without a standardized mechanism to invoke such external functions. With MCP, a protocol layer is provided that allows external functions to be described, discovered, and invoked at runtime – without requiring prior hard-coded integration into the LLM.
MCP follows a client-server architecture: MCP servers expose functional interfaces, metadata, and invocation mechanisms; MCP clients, e.g., LLM-based agents, can dynamically query these servers to discover available operations and execute them by issuing well-defined requests. [Ant_ModelContextProtocolSpecification_2025]
@@ III Related Work
A recent survey in [Ray_ASurveyonModel_2025] outlines MCP ’s core concepts and its potential to expose external tools to LLMs in a standardized way. The authors discuss possible applications of MCP in domains such as manufacturing, where LLM agents might retrieve sensor data, interact with Manufacturing Execution System (MES) platforms, or initiate control actions in response to anomalies. However, Ref. [Ray_ASurveyonModel_2025] highlights only conceptual opportunities and does not present concrete implementations or evaluate the feasibility of using LLMs for tool selection or task execution in physical automation systems.
@@
While the authors of [Ray_ASurveyonModel_2025] outline the conceptual potential of the new MCP for exposing tools to LLMs, other recent studies focus more directly on how LLMs can be applied to planning and control tasks in industrial automation. Xia et al. present in [XJZ+_ControlIndustrialAutomationSystem_2024] an agent-based framework using structured prompts, event-driven semantics, and a hierarchy of manager and operator agents that interact with automation modules via digital twins. The control relies on generated function calls and standard operation procedures as a knowledge base, but requires fine-tuning, curated datasets, and tightly engineered prompt structures – limiting flexibility and generalizability.
@@
A similar direction is pursued in the LLMAPM framework introduced in [NWL+_Alargelanguagemodelbased_2025], which targets manufacturing process planning. Natural language task descriptions are decomposed into workflows, verified through state machines, and executed on heterogeneous devices via a low-code platform. While supporting human-in-the-loop corrections, the system still depends on handcrafted prompt templates and lacks a runtime interface for tool abstraction and reuse.
@@
In [VyMe_AutonomousIndustrialControlusing_2024] a multi-agent LLM control framework uses validator and reprompter agents with a digital twin to iteratively refine actions for robust decision-making in industrial settings. While this enables safe and adaptive control, it is demonstrated on a single closed-loop task and does not generalize to dynamic function composition or modular tool invocation.
@@
The use of LLMs in multi-robot systems is reviewed by Li et al. in [LZA+_LargeLanguageModelsfor_2025], covering applications in task allocation and coordination. While relevant for distributed systems, the focus remains on agent negotiation and motion planning rather than exploring how LLMs might interface with standardized protocols like MCP for structured tool invocation.
@@
A broader perspective on LLM-driven agent systems is provided in [LiTo_LLMPoweredAIAgentSystems_2025], which categorizes LLM-powered agents into software-based, physical, and adaptive hybrid types. While the survey highlights potential use cases in smart manufacturing, it acknowledges ongoing challenges such as output uncertainty, integration complexity, and lack of standardized tool access in physical environments.
@@ IV-A MCP Architecture
In the proposed architecture (see Figure 1), user interfaces and control systems like MES act as MCP clients that are directly connected to an LLM. MCPs servers provide functions, resources, and prompts. According to the definition in [KBH+_AReferenceModelfor_15.09.2022b], a capability is defined as a specification of a function with either physical or virtual effects. This distinction of effects is reflected in the MCP architecture with the two alternatives of direct MCP servers and gateway MCP servers.
@@
User queries or control requests are sent together with all available MCP servers as prompts to the LLM (1a). Depending on the response of the LLM (1b), MCP functionalities are selected and used by the MCP client. In case of direct MCP servers, this functionality is called (2a) and the returned result (2b) is interpreted. In the case of gateway MCP servers, the function call (3a) may lead to an external request (3b) that yields a return (3c), which is processed by the gateway MCP server and then returned to the MCP client (3d). The following subsection illustrates how elements of the CSS reference model presented in [KBH+_AReferenceModelfor_15.09.2022b] relate to elements in this MCP architecture.
@@ IV-B Representing CSS Elements in MCP
Each Capability is a function declaration in one of the programming frameworks provided by [Ant_ModelContextProtocolSpecification_2025]. Its specification details are documented in natural language using descriptions and code comments. This natural-language description is the counterpart of a capability model and enables LLMs to understand the purpose, constraints, and semantics of a capability.
@@
Constraints are subdivided into PropertyConstraints and TransitionConstraints [KBH+_AReferenceModelfor_15.09.2022b]. PropertyConstraints are internal to a capability. They can be expressed as comments and embedded as assertions or preconditions inside the capability function. Violations can raise structured errors, which the LLM can interpret (e.g., drill diameter $\leq 50$). TransitionConstraints are global constraints between multiple capabilities: These cannot be checked by one capability alone, but instead must be described textually (e.g., “may only be used after Capability X”) and are interpreted by the LLM during invocation.
A Skill is the implementation of a capability, i.e., the function definition. This is in line with the specification of a skill in [KBH+_AReferenceModelfor_15.09.2022b]. However, MCP provides an alternative interface to expose and describe the skill’s callable entry points.
@@ V Evaluation
    Standard Process Execution: A user specifies a workpiece to be drilled (i.e. defines material and diameter) and transported to an assembly station. The LLM correctly infers that the spindle speed must be calculated first. It uses the calculation skill to retrieve a revolutions per minute (RPM) value, invokes the drilling skill with this RPM value, and dispatches the mobile robot for transport. All skills were executed in correct order and with correct parameters.
@@ VI Conclusion and Future Work
This paper presents a lightweight and LLM-driven alternative to current model-based CSS approaches in manufacturing. By leveraging MCP, we demonstrate how capabilities can be described using comments and skills can be exposed as callable tools (RQ1). Through a set of scenarios, we demonstrate the potential of an LLM paired with MCP to plan and execute multi-step manufacturing processes, discover available capabilities, interpret their purposes, invoke corresponding skills, and handle errors or ambiguities gracefully – without explicit information models, e.g., ontologies or AASs. Our results suggest that explicit CSS models may become less necessary in the future, as LLMs increasingly infer intent from text and function implementations (RQ2). This would significantly lower the barrier to integration and reconfiguration in flexible manufacturing.
However, our results are only an initial validation of LLM-based orchestration via MCP, and reliability and robustness – including hallucinations – remain an open challenge with many directions for future work. Instead of a single LLM orchestrator, future approaches should use multiple, dedicated LLM-based agents with defined roles (e.g., planner, verifier, negotiator) for more scalable, modular, and robust architectures. In addition, decentralized negotiation patterns to compare service offers across organizational boundaries may be realized by LLM agents.
Another future research direction is to enable LLMs not only to select and invoke skills, but to also generate new ad-hoc skills on demand. Building on earlier work on skill generation [LAN+_CapabilityDriven_2025], a combined approach could allow LLMs to identify missing functionality, synthesize MCP tool code, register the new tool, and invoke it as part of a task plan. The approach of ”model-free” automation, where planning and execution emerge from linguistic reasoning rather than predefined symbolic models, can also be applied to simplify integration processes in other domains, e.g., modular electrolyzer systems.